# Multi-type branching processes with immigration generated by point processes

### v1

@@ 5 Proofs
For the first two items we use techniques from the works of Yanev et al. [15, 16, 42, 44], for which we will need the following lemma.
Lemma 5.1 (Lemmas 2 and 3 from [42]).
Let $\lambda_{\infty}$ and $\psi_{\infty}$ be positive real constants.
  1. 1.
    If $\lambda(t)\stackrel{{\scriptstyle\infty}}{{\sim}}\lambda_{\infty}e^{\delta t}$ with $\delta\in\mathbb{R}$ and $\int_{0}^{\infty}e^{-\delta x}\psi(x)\,\mathrm{d}x$ finite, then
    $\int_{0}^{t}\lambda(t-u)\psi(u)\mathrm{d}u\stackrel{{\scriptstyle\infty}}{{\sim}}\lambda(t)\int_{0}^{\infty}e^{-\delta x}\psi(x)\mathrm{d}x.$
  2. 2.
    If $\lambda(t)\stackrel{{\scriptstyle\infty}}{{\sim}}\lambda_{\infty}$ and $\psi(t)\stackrel{{\scriptstyle\infty}}{{\sim}}\psi_{\infty}$, then
    $\int_{0}^{t}\lambda(t-u)\psi(u)\mathrm{d}u\stackrel{{\scriptstyle\infty}}{{\sim}}\lambda_{\infty}\psi_{\infty}t.$
For items 3 and 4 we follow an approach similar to the one chosen by Rabehasaina and Woo in [35].
Let us start with the observation that the stationarity assumption (4.7), applied for $n=1$, implies that $K(x,x)$ is constant over $x$. Denote its value by $K_{\ast}$.
(1) We have by Theorem 4.3 that $\mathbb{E}\left[Z_{i}(t)\right]=K_{\ast}\int_{0}^{t}\mathbb{E}\left[Z_{\times,i}(t-x)\right]\lambda(x)\mathrm{d}x$. Furthermore, by classical results, see e.g. [40, p.151] or [2, p. 203],
$\mathbb{E}\left[Z_{\times,i}(t)\right]\stackrel{{\scriptstyle\infty}}{{\sim}}e^{\rho t}a_{i}:=e^{\rho t}\langle\boldsymbol{u},\mathbb{E}\left[\boldsymbol{I}\right]\rangle v_{i}.$
Therefore, applying the first part of Lemma 5.1 with $\psi\equiv\mathbb{E}\left[Z_{\times,i}(\cdot)\right]$, we obtain
$\mathbb{E}\left[Z_{i}(t)\right]\stackrel{{\scriptstyle\infty}}{{\sim}}e^{\delta t}A_{i}=:e^{\delta t}K_{\ast}\lambda_{\infty}a_{i}\int_{0}^{\infty}e^{-\delta x}\mathbb{E}\left[Z_{\times,i}(x)\right]\mathrm{d}x.$
We prove that $Z_{i}(t)$ concentrates around its mean by bounding its variance: by the second part of Theorem 4.3,
$Var(Z_{i}(t))\leq K_{\ast}\int_{0}^{t}\mathbb{E}\left[\left(Z_{\times,i}(t-x)\right)^{2}\right]\lambda(x)\mathrm{d}x.$
The asymptotics of the factorial moments of $Z_{\times,i}(t)$ are available in [40, p.152] or [2, p.204], giving us, for some positive constants $B_{j,i}$,
$\mathbb{E}\left[\left(Z_{\times,i}(t)\right)^{2}\right]\stackrel{{\scriptstyle\infty}}{{\sim}}\begin{cases}e^{\rho t}B_{1,i}&\text{if $\rho<0$},\\ tB_{2,i}&\text{if $\rho=0$},\\ e^{2\rho t}B_{3,i}&\text{if $\rho>0$}.\end{cases}$
Therefore, using again the first part of Lemma 5.1, and its second part for the case $2\rho=\delta$, we get that in the case where $\delta>\max\{0,\rho\}$,
$\quad\text{Var}(Z_{i}(t))=\begin{cases}O(e^{\delta t})&\text{if $\rho\leq 0$},\\ O(e^{\max\{2\rho,\delta\}t})&\text{if $\rho>0$ and $\delta\neq 2\rho$},\\ O(te^{\delta t})&\text{if $2\rho=\delta>0$}.\end{cases}$
Since we have already seen that $\mathbb{E}\left[Z_{i}(t)\right]\sim e^{\delta t}A_{i}$ and $\delta>\max\{0,\rho\}$ by assumption, we infer that always
$\mathbb{E}\left[\left(\frac{Z_{i}(t)}{\mathbb{E}\left[Z_{i}(t)\right]}-1\right)^{2}\right]=\frac{Var(Z_{i}(t))}{\mathbb{E}\left[Z_{i}(t)\right]^{2}}\xrightarrow[t\to\infty]{}0$ (5.9)
as claimed.
(2) Similarly to (1), Using the previously referred results of Sevastyanov, and Athreya and Ney, we have that, as $\delta=\rho>0$,
$\mathbb{E}\left[Z_{\times,i}(t)\right]\stackrel{{\scriptstyle\infty}}{{\sim}}e^{\rho t}a_{i},\quad\text{and}\quad\mathbb{E}\left[(Z_{\times,i}(t))^{2}\right]\stackrel{{\scriptstyle\infty}}{{\sim}}e^{2\rho t}\left(B_{3,i}-a_{i}^{2}\right).$
Using again Lemma 5.1, we get that
$\mathbb{E}\left[Z_{i}(t)\right]=K_{\ast}\int_{0}^{t}\mathbb{E}\left[Z_{\times,i}(t-x)\right]\lambda(x)\mathrm{d}x\sim te^{\rho t}A^{\prime}_{i}:=te^{\rho t}K_{\ast}\lambda_{\infty}a_{i},\quad\text{and}\quad Var(Z_{i}(t))=\mathrm{O}\left(e^{2\rho t}\right).$
Therefore, the argument from (5.9) is applicable which establishes the result.
(4) In this case we assume that the process is subcritical ($\rho<0$) and that $\lambda(x)\sim\lambda_{\infty}$. By assumption
@@
$\left|1-\mathcal{L}_{\boldsymbol{Z}_{\times}}(x,\boldsymbol{s})\right|=\mathbb{E}\left[1-\exp\left(\langle\boldsymbol{Z}_{\times}(x),\boldsymbol{s}\rangle\right)\right]\leq\mathbb{E}\left[\langle\boldsymbol{Z}_{\times}(x),\boldsymbol{s}\rangle\right]\leq e^{\rho x}C_{\boldsymbol{s}}.$ (5.13)
@@
The last is an $\mathrm{d}\boldsymbol{x}$-integrable function with
$\begin{split}\int_{(0,\infty)^{n}}K_{\ast}^{n}C_{\boldsymbol{s}}^{n}&\prod_{i=1}^{n}\left(e^{\rho x_{i}}\sup_{x\geq C}\lambda(x)+\mathbbm{1}_{\left\{x_{i}\in(0,C)\right\}}\lambda(x_{i})\right)\mathrm{d}\boldsymbol{x}\\ &=\left(K_{\ast}C_{\boldsymbol{s}}\int_{0}^{\infty}\left(e^{\rho x}\sup_{x\geq C}\lambda(x)+\mathbbm{1}_{\left\{x_{i}\in(0,C)\right\}}\lambda(x)\right)\mathrm{d}x\right)^{n}<\infty,\end{split}$ (5.14)
because by assumption the directing measure of a DPP $\Lambda$ is always $\sigma$-finite and because in this case $\rho<0$. Therefore by dominated convergence
@@
$t\to\infty]{}\lambda_{\infty}^{n}\int_{(0,\infty)^{n}}D(\boldsymbol{x})\prod_{i=1}^{n}\left(1-\mathcal{L}_{\boldsymbol{Z}_{\times}}(x_{i},\boldsymbol{s})\right)\mathrm{d}\boldsymbol{x},$
@@
To conclude that the last is the Laplace transform of a positive random variable, it remains to verify that it is continuous at one, see e.g. [11, p. 431, Chapter XIII.1, Theorem 2]. This is indeed true by the bound (5.13) and the choice $C_{\boldsymbol{s}}=\max_{j}{s_{j}}/\min_{j}{v_{j}}$.
(3) For critical processes without immigration, the limit distribution is obtained under optimal assumptions (and for more general branching mechanisms) by Holte [14]. Again, see Remark 4.6, one should take into account the difference between the definition of $\boldsymbol{v}$ in [14] and in this work, notably that there are considered the eigenvectors of $\mathbb{E}\left[(\nu_{i})_{j}\right]$, compare with (4.4). Denoting with upper index $H$ the corresponding quantities from the work of Holte, and taking into account the normalization there $\langle\boldsymbol{u^{H}},\boldsymbol{v^{H}}\rangle=1,\langle\boldsymbol{v^{H}},\boldsymbol{1}\rangle=1$, we have that
$u_{i}^{H}=\frac{\sum_{j}(v_{j}/\mu_{j})}{\sum_{j}(u_{j}v_{j}/\mu_{j})}u_{i},\quad v_{i}^{H}=\frac{1}{\sum_{j}(v_{j}/\mu_{j})}\frac{v_{i}}{\mu_{i}},\quad\text{and}\quad\sum_{j}\mu_{j}u_{j}^{H}v_{j}^{H}=\frac{1}{\sum_{j}(u_{j}v_{j}/\mu_{j})}.$
Therefore from [14, Lemma on page p.493], we have that
$t\left(1-\mathcal{L}_{\boldsymbol{Z}_{\times}}\left(t,\frac{\boldsymbol{s}}{t}\right)\right)\xrightarrow[t\to\infty]{}{\beta}\frac{\langle\boldsymbol{s},\boldsymbol{v}\rangle}{Q+\boldsymbol{\langle s,\boldsymbol{v}\rangle}},$
where
$\beta:=\frac{\mathbb{E}\left[\langle\boldsymbol{I},\boldsymbol{u}\rangle\right]}{Q}\quad\text{with}\quad Q:=\frac{1}{2}\sum_{i,j,k=1}^{d}\frac{\partial^{2}G_{\boldsymbol{\nu_{i}}}}{\partial x_{j}\partial x_{k}}{v}_{i}{u}_{j}{u}_{k}.$
We will use the last limit in a form which is obtained by substituting, for $x\in(0,1)$ $t$ with $t(1-x)$, and $\boldsymbol{s}$ with $\boldsymbol{s}(1-x)$, which gives
$t\left(1-\mathcal{L}_{\boldsymbol{Z}_{\times}}\left(t(1-x),\frac{\boldsymbol{s}}{t}\right)\right)\xrightarrow[t\to\infty]{}\beta\frac{\langle\boldsymbol{s,\boldsymbol{v}}\rangle}{Q+(1-x)\boldsymbol{\langle s,\boldsymbol{v}\rangle}}.$
Substituting in (5.10) and changing variables $x_{i}\to tx_{i}$, we get
@@
Next, (5.11), (5.13) gives constant bounds of $D(t\boldsymbol{x})$ and the product. Moreover, as $\lambda$ is integrable, it is bounded for $||x||>\epsilon$. We also have that $D(t\boldsymbol{x})$ tends to 1 $\mathrm{d}\boldsymbol{x}$-a.e., because for $\boldsymbol{x}$ with different entries, $K(tx_{i},tx_{j})=K(t(x_{i}-x_{j}),0)$ tends to 0 if $x_{i}\neq x_{j}$ for $x\to\infty$ by assumption. Therefore by dominated convergence,
$\mathcal{L}_{\boldsymbol{Z}}\left(t,\frac{\boldsymbol{s}}{t}\right)$ $t\to\infty]{}1+\sum_{n\geq 1}\frac{(-1)^{n}}{n!}\left(\int_{0}^{1}\lambda_{\infty}\beta\frac{\langle\boldsymbol{s},\boldsymbol{v}\rangle}{Q+(1-x)\boldsymbol{\langle s,\boldsymbol{v}\rangle}}\mathrm{d}x\right)^{n}$
$=\left(\frac{Q}{Q+\langle\boldsymbol{s},\boldsymbol{v}\rangle}\right)^{\lambda_{\infty}\beta}.$
The last is indeed the Laplace transform of $Y\boldsymbol{v}$ with $Y\sim Gamma(\lambda_{\infty}\beta,Q)$ as claimed, which concludes the proof.

### v2

@@ 6 Proofs
For items 3 and 4 we follow an approach similar to the one chosen by Rabehasaina and Woo in [36], and for 1 and 2, techniques from the works [15, 16, 44, 46], which core argument lies in Lemma 5.5.
Indeed substituting its estimates, we see that if $\delta>\max\{\rho,0\}$ or $\delta=\rho>0$,
$\mathbb{E}\left[\left(\frac{Z_{i}(t)}{\mathbb{E}\left[Z_{i}(t)\right]}-1\right)^{2}\right]=\frac{\operatorname{Var}(Z_{i}(t))}{\mathbb{E}\left[Z_{i}(t)\right]^{2}}\xrightarrow[t\to\infty]{}0,$ (6.11)
which gives exactly items 1 and 2 of the theorem.
(4) In this case we assume that the process is subcritical ($\rho<0$) and that $\lambda(x)\stackrel{{\scriptstyle\infty}}{{\sim}}\lambda_{\infty}$. As we assume that the immigration is spanned by a DPP, by Theorem 4.2
@@
$\left|1-\mathcal{L}_{\boldsymbol{Z}_{\times}}(x,\boldsymbol{s})\right|=\mathbb{E}\left[1-\exp\left(-\langle\boldsymbol{Z}_{\times}(x),\boldsymbol{s}\rangle\right)\right]\leq\mathbb{E}\left[\langle\boldsymbol{Z}_{\times}(x),\boldsymbol{s}\rangle\right]\leq e^{\rho x}C_{\boldsymbol{s}}.$ (6.15)
@@
The last is a $\mathrm{d}\boldsymbol{x}$-integrable function with
$\begin{split}\int_{(0,\infty)^{n}}K_{\ast}^{n}C_{\boldsymbol{s}}^{n}&\prod_{i=1}^{n}\left(e^{\rho x_{i}}\sup_{x\geq C}\lambda(x)+\mathbbm{1}_{\left\{x_{i}\in(0,C)\right\}}\lambda(x_{i})\right)\mathrm{d}\boldsymbol{x}\\ &=\left(K_{\ast}C_{\boldsymbol{s}}\int_{0}^{\infty}\left(e^{\rho x_{i}}\sup_{x\geq C}\lambda(x)+\mathbbm{1}_{\left\{x_{i}\in(0,C)\right\}}\lambda(x)\right)\mathrm{d}x\right)^{n}<\infty,\end{split}$ (6.16)
because by assumption the directing measure $\Lambda$ of a DPP is always $\sigma$-finite and because in this case $\rho<0$. Therefore by dominated convergence
@@
$\xrightarrow[t\to\infty]{}\lambda_{\infty}^{n}\int_{(0,\infty)^{n}}D(\boldsymbol{x})\prod_{i=1}^{n}\left(1-\mathcal{L}_{\boldsymbol{Z}_{\times}}(x_{i},\boldsymbol{s})\right)\mathrm{d}\boldsymbol{x},$
@@
To conclude that the last is the Laplace transform of a positive random variable, it remains to verify tends to $1$ as $\boldsymbol{s}\to\boldsymbol{0}$, see e.g. [11, p. 431, Chapter XIII.1, Theorem 2]. This is indeed true by the bound (6.15) and the choice $C_{\boldsymbol{s}}=\max_{j}{s_{j}}/\min_{j}{v_{j}}$.
(3) For critical processes without immigration, Holte [14] obtains the limit distribution under optimal assumptions and for a more general class of branching mechanisms. Note, however, that Holte’s $\boldsymbol{v}$ is an eigenvector of the matrix with entries $\mathbb{E}\left[(\nu_{i})_{j}\right],$ whereas in this work we define $\boldsymbol{v}$ via (4.4) (see also Remark 5.4). Denoting with lower index $H$ the corresponding quantities from the work of Holte, and taking into account the normalization there $\langle\boldsymbol{u}_{H},\boldsymbol{v}_{H}\rangle=1,\langle\boldsymbol{v}_{H},\boldsymbol{1}\rangle=1$, we have that for $D_{1}:={\sum_{j}(v_{j}/\mu_{j})}$ and $D_{2}:=\sum_{j}(u_{j}v_{j}/\mu_{j})$,
$u_{H,i}=\frac{D_{1}}{D_{2}}u_{i},\quad v_{H,i}=\frac{1}{D_{1}}\frac{v_{i}}{\mu_{i}},\quad\text{and}\quad\sum_{j}\mu_{j}u_{H,j}v_{H,j}=\frac{1}{D_{2}}.$ (6.18)
We show how the results of Holte [14] translate to our setting. Let $(e_{i})$ be the standard basis in $\mathbb{R}^{d}$. Therefore from [14, Lemma on page p.493], and using lower indices $H$ for constants from [14], we have that
$H_{i}\left(\frac{\boldsymbol{s}}{t},t\right)$ $:=\mathbb{E}\left[1-\exp\left.\left(-\langle Z_{\times}(t),\frac{\boldsymbol{s}}{t}\rangle\right)\right|Z_{\times}(0)=e_{i})\right]$
[…]
Our process starts with a random number of particles of law $\boldsymbol{I}$, so a first order expansion gives us
$\begin{split}1-\mathcal{L}_{\boldsymbol{Z}_{\times}}\left(t,\frac{\boldsymbol{s}}{t}\right)&=\mathbb{E}\left[1-\exp\left(-\langle\boldsymbol{Z}_{\times}(t),\frac{\boldsymbol{s}}{t}\rangle\right)\right]\\ &=\sum_{\boldsymbol{x}=(x_{1},\dots,x_{d})}\mathbb{P}(\boldsymbol{I}=\boldsymbol{x})\left(1-\prod_{i}{\left(1-H_{i}\left(\frac{\boldsymbol{s}}{t},t\right)\right)}^{x_{i}}\right)\\ &=\sum_{\boldsymbol{x}=(x_{1},\dots,x_{d})}\mathbb{P}(\boldsymbol{I}=\boldsymbol{x})\left(1-\prod_{i}\left(1-x_{i}\frac{C_{i}}{t}+\mathrm{o}\left(\frac{1}{t}\right)\right)\right)\\ &=\sum_{\boldsymbol{x}=(x_{1},\dots,x_{d})}\mathbb{P}(\boldsymbol{I}=\boldsymbol{x})\left(\sum_{i}\left(x_{i}\frac{C_{i}}{t}+\mathrm{o}\left(\frac{1}{t}\right)\right)\right)\\ &=\frac{1}{t}\mathbb{E}\left[\langle\boldsymbol{I},{\boldsymbol{C}}\rangle\right]+\mathrm{o}\left(\frac{1}{t}\right),\quad\text{with }\boldsymbol{C}:=(C_{1},\dots,C_{d}).\end{split}$ (6.19)
Setting our parameters
$\beta:=\frac{\mathbb{E}\left[\langle\boldsymbol{I},\boldsymbol{u}\rangle\right]}{Q}\quad\text{with}\quad Q:=\frac{1}{2}\sum_{i,j,k=1}^{d}\left.\frac{\partial^{2}G_{{\nu_{i}}}}{\partial x_{j}\partial x_{k}}\right|_{\boldsymbol{x}=\boldsymbol{1}}\mu_{i}^{-1}{v}_{i}{u}_{j}{u}_{k},$
using (6.18), we calculate the constants from [14]:
[…]
Substituting the definition of $C_{i}$ in (6.19), we obtain
$t\left(1-\mathcal{L}_{\boldsymbol{Z}_{\times}}\left(t,\frac{\boldsymbol{s}}{t}\right)\right)\xrightarrow[t\to\infty]{}{\beta}\frac{\langle\boldsymbol{s},\boldsymbol{v}\rangle}{1/Q+\boldsymbol{\langle s,\boldsymbol{v}\rangle}}.$
[…]
We will use the last limit in a form which is obtained by substituting $t$ with t(1x)t(1-x), and $\boldsymbol{s}$ with $\boldsymbol{s}(1-x)$ for $x\in(0,1)$, which gives
$t(1-x)\left(1-\mathcal{L}_{\boldsymbol{Z}_{\times}}\left(t(1-x),\frac{\boldsymbol{s}}{t}\right)\right)\xrightarrow[t\to\infty]{}\beta\frac{(1-x)\langle\boldsymbol{s,\boldsymbol{v}}\rangle}{1/Q+(1-x)\boldsymbol{\langle s,\boldsymbol{v}\rangle}}.$
Dividing by (1x)(1-x), then substituting in (6.12) and changing variables $x_{i}\to tx_{i}$, we get
@@
Next, (6.13) and (6.15) give constant bounds of $D(t\boldsymbol{x})$ and the product. Moreover, as $\lambda$ is integrable, it is bounded on $|x|>\epsilon$. We also have that $D(t\boldsymbol{x})$ tends to 1 $\mathrm{d}\boldsymbol{x}$-a.e., because for $\boldsymbol{x}$ with different entries, $K(tx_{i},tx_{j})=K(t|x_{i}-x_{j}|,0)$ tends to 0 if $x_{i}\neq x_{j}$ for $t\to\infty$ by assumption. Therefore by dominated convergence,
$\mathcal{L}_{\boldsymbol{Z}}\left(t,\frac{\boldsymbol{s}}{t}\right)$ $\xrightarrow[t\to\infty]{}1+\sum_{n\geq 1}\frac{(-1)^{n}}{n!}\left(K^{\ast}\int_{0}^{1}\lambda_{\infty}\beta\frac{\langle\boldsymbol{s},\boldsymbol{v}\rangle}{1/Q+(1-x)\boldsymbol{\langle s,\boldsymbol{v}\rangle}}\mathrm{d}x\right)^{n}$
$=\left(\frac{1/Q}{1/Q+\langle\boldsymbol{s},\boldsymbol{v}\rangle}\right)^{K^{\ast}\lambda_{\infty}\beta}.$
The last is indeed the Laplace transform of $Y\boldsymbol{v}$ with $Y\sim\Gamma(K^{\ast}\lambda_{\infty}\beta,1/Q)$ as claimed, which concludes the proof.
@@ Acknowledgments
We gratefully thank the anonymous referees for their careful review of our manuscript and for their valuable comments and suggestions.